OpenStudy (anonymous):

A cheetah ran 300 feet in 2.92 seconds. What was the cheetah’s average speed in miles per hour?

OpenStudy (anonymous):

convert seconds to hour and feet to miles

OpenStudy (baldymcgee6):



OpenStudy (tkhunny):

\[\frac{300\;ft}{2.92\;sec}\cdot\frac{1\;mile}{5280\;ft}\cdot\frac{60\;sec}{1\;min}\cdot\frac{60\;min}{1\;hr}\] It's a while lot of multiplying by one (1).